# DvD: Unleashing a Generative Paradigm for Document Dewarping via Coordinates-based Diffusion Model

### v1

@@ 4.1. AnyPhotoDoc6300 Benchmark
Despite significant advances in document dewarping, the development of corresponding benchmarks lag behind. We summarized current benchmarks in Tab. 1, and we can find most of datasets suffer from restricted coverage of scenarios, small size of dataset, and deficient annotations of domains, impeding a comprehensive evaluation of dewarping models. To this end, we build a new large-scale benchmark AnyPhotoDoc6300, containing 6,300 real-world photographic document pairs. In our AnyPhotoDoc6300, we provide three distinct domain annotations to enable a more fine-grained quantitative evaluation, including layout category (LC), environment lighting (EL), and capture angles (CA). In addition, we aim to evaluate the dewarping capability of the model on three typical warping patterns (i.e., curves, folds, and crumples) under any combination of three types of given domains. Fig. 4 visualizes a sample array of ”Complex layout paper” (one of the eight layout categories). By meticulously specifying three environmental lighting and two capture angles, we can form any domain combination for the same document content. Consequently, we can obtain a fine-grained performance evaluation like Fig. 5 to discover the underlying issues for the current dewarping model. In Appendix E, we provide more details about our AnyPhotoDoc6300 benchmark, including collection settings and more visualization from different layout categories. We put some examples of AnyPhotoDoc6300 benchmark in supplementary materials and the completed benchmark will be publicly available after the paper review.
@@ 4.2. Metrics
Feature Alignment metrics. Following most of previous models (Ma et al., 2018; Das et al., 2019; Ma et al., 2022), we adopt the MS-SSIM (Multi-scale Structural Similarity) (Wang et al., 2003), LD (Local Distortion) (You et al., 2017b), and AD (Aligned Distortion) (Ma et al., 2022) to evaluate differences between dewarped and flat ground truth. Among them, MS-SSIM focuses on perceiving similarity on luminance, contrast, and structural information. While LD and AD focus on measuring the variations of SIFT flow (Liu et al., 2010).
OCR metrics. To evaluate the OCR readability improvement enabled by document dewarping, measuring recognition discrepancy between dewarped and flat documents has become the de facto standard in contemporary document dewarping models. Concretely, an off-the-shelf OCR engine is applied to recognize two text sequences from dewarped and flat documents, respectively. Then, ED (Edit Distance) and CER (Character Error Rate)  (Levenshtein et al., 1966) are harnessed to quantify the degree of deviation between two sequences. Pioneering a metric supplement, we further extend current OCR metrics in document dewarping by replacing previous OCR engines with MLLMs. Witnessing the superior progress of MLLMs in OCR capabilities recently (Wei et al., 2024; Karmanov et al., 2025; Nassar et al., 2025), we identify that there is still no exploration about whether the dewarped document can attain equivalent readability to its flat counterpart for MLLMs. To fill the blank, we pioneer MLLM-based OCR metrics (i.e., MMCER, MMED) to serve as a specialized supplement for prevalent ED and CER. Specifically, given the prompt of ”OCR the plain text” as a fixed instruction, we employ open-source MLLM QWen2.5-VL 7B (Bai et al., 2025) to recognize all characters in both dewarped and flat documents for ED and CER calculation.
@@ 4.3. Qualitative and Quantitative Comparison
Qualitative Comparisons. Dewarped document results on the AnyPhotoDoc6300 and DocUNet benchmark are shown in Fig. 6 and Figure-only pages. In Fig. 6, we select the five most recent open‑source dewarping models as well as GPT‑4o’s native generation model. On figure‑only pages, we additionally compare against publicly released inference results from non‑open‑source methods. Basically, our DvD achieves precise structure preservation in both local and overall document content. In contrast, GPT‑4o tends to produce unfaithful results that look visually clean but whose content is often chaotic. We argue this is because the image translation paradigm adopted by GPT-4o lacks explicit deformation awareness.
@@
Quantitative Comparisons. We compare the performance of our DvD model with previous state-of-the-art on three benchmarks, including DocUNet (Ma et al., 2018), DIR300 (Feng et al., 2022), and the proposed AnyPhotoDoc6300. The quantitative results on the DocUNet benchmark are shown in Tab. 2. For DIR300, we have placed the results in Appendix C. The ”Warped Document” in the first row means that we simply feed the raw input image for evaluation, therefore, most of the metrics perform the worst. Since our paper focuses on novel model designs rather than contributing high-quality data, we only train DvD on the currently most widely used Doc3D (Das et al., 2019) dataset. For a fair comparison, we divide the current methods into two branches according to their training data. It can be seen that even though we only used the Doc3D dataset, our method still achieved the best performance on the majority of metrics. In the upper first branch, DvD can achieve a slight overtaking, while in the lower second branch, DvD can achieve a significant superiority under a uniform Doc3D dataset. To be noted, compared with the first row, our experiments on MMCER and MMED reveal counterintuitive performance degradation in earlier dewarping models (e.g., DispFlow, DewarpNet), which we attribute to the fact that MLLMs are sensitive to resolution reduction and artifact (Li et al., 2024; Feng et al., 2024). On the other hand, compared with the first row, our DvD significantly reduces MMCER and MMED, and greatly improves the perception of photographed documents for MLLMs.
@@
Fine-grained Quantitative Comparisons. Fig. 5 illustrates AD, MS‑SSIM, and MMCER as three representative metrics via the Any-PhotoDoc6300 benchmark. More metrics in different domains are also exhibited in the Figures-only pages. Our DvD comprehensively achieves superior performance across various domains, including layout, lighting, and angles. Moreover, for the first time, we can also unveil performance of dewarping models at a fine-grained level. For instance, In Fig. 5 (I), existing methods generally suffer from poor AD performance on the warping pattern of the curves under moderate or warm lighting. In Fig. 5 (II), we observe a notable MS-SSIM drop for dark lighting documents with folds and crumples. In Fig. 5 (III), warm lighting causes the most pronounced decline in OCR metrics, especially on documents with curves. By pinpointing these issues unconventionally, we expect to provide the research community deeper insights into dewarping model behaviors and dataset curation, for driving further performance improvements.
@@ 4.4. Ablation Studies
Effectiveness on different Dewarping Paradigms. To verify the superiority of the proposed paradigm with the regression-based paradigm, we specially train another network of DvD by directly regressing the mapping. Then we can fairly compare different learning paradigms under the same network structure. As demonstrated in Tab. 3, The DvD baseline model trained using the regression-based paradigm has led to a general decline in performance, which emphasizes the effectiveness of our mapping generation paradigm for obtaining a more precise structure preservation.
Component Analysis of Compound Conditions. Tab. 4 demonstrates three types of compound condition $c_{t}$ configurations. We cam see that only using raw document feature $f_{d}$ does not obtain satisfied performance, which is then improved by adding document foreground $f_{m}$ and text-lines $f_{l}$. Finally, adding time-variant condition $r_{t}$ further boosts the performance obviously on all metrics. All of these verify the complementarity of those compound conditions.
@@
Table 5. Ablations on diffusion step, performance, and time consumption.
@@
Computational Efficiency Analysis. Tab. 5 illustrates a trade-off between performance and time consumption, driven by different diffusion denoising steps. As the number of steps increases from 1 to 3, the model’s performance improves substantially. However, beyond 3 steps, performance largely plateaus or even slightly declines. At the same time, increasing the step count incurs a linearly growing time cost. In our experiments, we set 3 as the optimal step setting for a trade off between performance and time consumption.
@@ 5. Conclusion
This paper unleashes a novel mapping generation paradigm for the document dewarping task by reformulating it as a coordinate-based denoising diffusion framework. To the best of our knowledge, this is the first attempt to explore the viability of dewarping a document using the diffusion model, where we tailor a coordinate‑level denoising strategy and a time-variant condition refinement (TVCR) mechanism, enabling precise preservation of document structures. To foster a fine-grained evaluation of dewarping models, we also build a new photographic document dewarping benchmark AnyPhotoDoc 6300, which is large-scale size, covers multiple scenarios and provides detailed domain annotations. Findings and insights from our experiments are poised to substantially advance photographic document processing and further impact a broad spectrum of graphics applications.

### v2

@@ 4.1. AnyPhotoDoc6300 Benchmark
Despite significant advances in document dewarping, the development of corresponding benchmarks lags behind. We summarized current benchmarks in Tab. 1, and we can find most of datasets suffer from restricted coverage of scenarios, small size of dataset, and deficient annotations of domains, impeding a comprehensive evaluation of dewarping models. To this end, we build a new large-scale benchmark AnyPhotoDoc6300, containing 6,300 real-world photographic document pairs. In our AnyPhotoDoc6300, we provide three distinct domain annotations to enable a more fine-grained quantitative evaluation, including layout category (LC), environment lighting (EL), and capture angles (CA). In addition, we aim to evaluate the dewarping capability of the model on three typical warping patterns (i.e., curves, folds, and crumples) under any combination of three types of given domains. Fig. 5 visualizes a sample array of ”Complex layout paper” (one of the eight layout categories). By meticulously specifying three environmental lighting and two capture angles, we can form any domain combination for the same document content. Consequently, we can obtain a fine-grained performance evaluation like Fig. 7 to discover the underlying issues for the current dewarping model. In Appendix E, we provide more details about our AnyPhotoDoc6300 benchmark, including data collection settings and more visualizations from different layout categories.
@@ 4.2. Metrics
4.2.1. Feature Alignment Metrics.
Following most of previous models (Ma et al., 2018; Das et al., 2019; Ma et al., 2022), we adopt the MS-SSIM (Multi-scale Structural Similarity) (Wang et al., 2003), LD (Local Distortion) (You et al., 2017b), and AD (Aligned Distortion) (Ma et al., 2022) to evaluate differences between dewarped and flat ground truth. Among them, MS-SSIM focuses on perceiving similarity on luminance, contrast, and structural information. While LD and AD focus on measuring the variations of SIFT flow (Liu et al., 2010).
Figure 7. Quantitative dewarping performance comparisons on the AnyPhotoDoc6600 benchmark dataset. Under the multiple Layout Pattern, we provided 18 dimensions of evaluation for AD, MS-SSIM, and MMCER. More evaluation results can be found on the Figures-only pages after the reference.
4.2.2. OCR Metrics.
To evaluate the OCR readability improvement enabled by document dewarping, measuring recognition discrepancy between dewarped and flat documents has become the de facto standard in contemporary document dewarping models. Concretely, an off-the-shelf OCR engine is applied to recognize two text sequences from dewarped and flat documents, respectively. Then, ED (Edit Distance) and CER (Character Error Rate)  (Levenshtein et al., 1966) are harnessed to quantify the degree of deviation between two sequences. Pioneering a metric supplement, we further extend current OCR metrics in document dewarping by replacing previous OCR engines with MLLMs. Witnessing the superior progress of MLLMs in OCR capabilities recently (Wei et al., 2024; Karmanov et al., 2025; Nassar et al., 2025), we identify that there is still no exploration about whether the dewarped document can attain equivalent readability to its flat counterpart for MLLMs. To fill the blank, we pioneer MLLM-based OCR metrics (i.e., MMCER, MMED) to serve as a specialized supplement for prevalent ED and CER. Specifically, given the prompt of ”OCR the plain text” as a fixed instruction, we employ open-source MLLM QWen2.5-VL 7B (Bai et al., 2025) to recognize all characters in both dewarped and flat documents for ED and CER calculation.
@@ 4.3. Qualitative and Quantitative Comparison
Dewarped document results on the AnyPhotoDoc6300 and DocUNet benchmark are shown in Fig. 6. In this figure, we select the five most recent open‑source dewarping models as well as GPT‑4o’s native generation model. Our prompt fed to GPT-4o is consistently given, i.e., ”Please perform dewarping on this document to make it flat and clear.”. On Fig. 8 and Fig. 9, we additionally compare against publicly released inference results from non‑open‑source methods. Basically, our DvD achieves precise structure preservation in both local and overall document content. In contrast, GPT‑4o tends to produce unfaithful results that look visually clean but whose content is often chaotic. We argue this is because the image translation paradigm adopted by GPT-4o lacks explicit deformation awareness.
4.3.2. Quantitative Comparisons.
@@
We compare the performance of our DvD model with previous state-of-the-art on three benchmarks, including DocUNet (Ma et al., 2018), DIR300 (Feng et al., 2022), and the proposed AnyPhotoDoc6300. The quantitative results on the DocUNet benchmark are shown in Tab. 2. For DIR300, we have placed the results in Appendix C. The ”Warped Document” in the first row means that we simply feed the raw input image for evaluation, therefore, most of the metrics perform the worst. Since our paper focuses on novel model designs rather than contributing high-quality data, we only train DvD on the currently most widely used Doc3D (Das et al., 2019) dataset. For a fair comparison, we divide the current methods into two branches according to their training data. It can be seen that even though we only used the Doc3D dataset, our method still achieved the best performance on the majority of metrics. In the upper first branch, DvD can achieve a slight overtaking, while in the lower second branch, DvD can achieve a significant superiority under a uniform Doc3D dataset. To be noted, compared with the first row, our experiments on MMCER and MMED reveal counterintuitive performance degradation in earlier dewarping models (e.g., DispFlow, DewarpNet), which we attribute to the fact that MLLMs are sensitive to resolution reduction and artifact (Li et al., 2024; Feng et al., 2024). On the other hand, compared with the first row, our DvD significantly reduces MMCER and MMED, and greatly improves the perception of photographed documents for MLLMs.
4.3.3. Fine-grained Quantitative Comparisons.
@@
Fig. 7 illustrates AD, MS‑SSIM, and MMCER as three representative metrics via the Any-PhotoDoc6300 benchmark. More metrics in different domains are also exhibited in the Fig. 10,11,12, and Appendix. Our DvD comprehensively achieves superior performance across various domains, including layout, lighting, and angles. Moreover, for the first time, we can also unveil some brand-new performance comparisons for dewarping models at a fine-grained level. For instance, in Fig. 7 (I), existing methods generally suffer from severe AD performance decline under mixed warm lighting and slant angle, especially on the warping pattern of the curves and crumple. In Fig. 7 (II), we observe a notable MS-SSIM drop for dark lighting documents. In Fig. 7 (III), warm lighting causes the most pronounced decline in OCR metrics, especially on crumpled documents. By pinpointing these issues unconventionally, we expect to provide the research community deeper insights into dewarping model behaviors and dataset curation, for driving further performance improvements.
@@ 4.4. Ablation Studies
4.4.1. Effectiveness on different Dewarping Paradigms.
To verify the superiority of the proposed paradigm over the regression-based paradigm, we specially train another network of DvD by directly regressing the mapping. Then we can fairly compare different learning paradigms under the same network structure. As demonstrated in Tab. 3, The DvD baseline model trained using the regression-based paradigm has led to a general decline in performance, which emphasizes the effectiveness of our mapping generation paradigm for obtaining a more precise structure preservation.
4.4.2. Component Analysis of Compound Conditions.
Tab. 4 demonstrates three types of compound condition $c_{t}$ configurations. We can see that only using the raw document feature $f_{d}$ does not obtain satisfactory performance, which is then improved by adding the document foreground $f_{m}$ and text-lines $f_{l}$. Finally, adding a time-variant condition $r_{t}$ further boosts the performance, obviously on all metrics. All of these verify the complementarity of those compound conditions.
@@
Table 5. Ablations on sampling step, performance, and time consumption.
@@
Table 6. Ablations study for different latent size.
[…]
$16\times 16$ 0.432 10.63 0.462 0.451 1343.87
$32\times 32$ 0.492 7.69 0.370 0.385 1023.28
$64\times 64$ 0.549 6.61 0.279 0.366 928.94
$128\times 128$ 0.551 6.62 0.276 0.376 940.43
4.4.3. Computational Efficiency Analysis.
Tab. 5 illustrates a trade-off between performance and time consumption, driven by different diffusion denoising steps. The time unit here refers to the average elapsed seconds per image we take to infer the DocUNet (Ma et al., 2018) benchmark. As the number of sampling steps increases from 1 to 3, the model’s performance improves substantially. However, beyond 3 steps, performance largely plateaus or even slightly declines. We consider that this might be due to excessive steps that could accumulate errors over time, causing generated samples to deviate from the real distribution. At the same time, increasing the step count incurs a linearly growing time cost. In our experiments, we set 3 as the optimal step setting for a trade-off between performance and time consumption.
4.4.4. Size Analysis for Latent Space.
Tab. 6 presents the performance of the DvD model under different latent space resolutions. When the resolution is too small ($16\times 16$), the model performs poorly. We attribute this to the fact that such a limited latent space sacrifices overmuch warping semantics, making it difficult to accurately represent the high-resolution (i.e., $2000\times 3000$) backward mapping $M_{0}$. Empirically, we find that a moderate resolution of $64\times 64$ is sufficient to provide warping semantics. Further increasing the resolution to $128\times 128$ yields no significant performance gain.
4.4.5. Limitations.
Our method still has two limitations. (1) Slow training: Unlike directly selecting the denoising time-step in vanilla DDIM, training with the TVCR mechanism requires sampling a few steps per iteration, causing a slower training speed. (2) Limited Generalization on unseen document types: Diffusion models excel at generating samples that conform to the training data distribution. Since the training set Doc3D (Das et al., 2019) contains many academic papers and magazines, DvD shows superiority on these seen document types (cf. Fig. 10,11,12). However, DvD’s improvements on unseen types (Invoices/Education scripts) are negligible (cf. Fig. 15,16 in Appendix).
@@ 5. Conclusion
This paper unleashes a novel mapping generation paradigm for the document dewarping task by reformulating it as a coordinate-based denoising diffusion framework. To the best of our knowledge, this is the first attempt to explore the viability of dewarping a document using the diffusion model, where we tailor a coordinate‑level denoising strategy and a time-variant condition refinement (TVCR) mechanism, enabling precise preservation of document structures. To foster a fine-grained evaluation of dewarping models, we also build a new photographic document dewarping benchmark, AnyPhotoDoc 6300, which is large-scale in size, covers multiple scenarios, and provides detailed domain annotations. Findings and insights from our experiments are poised to substantially advance photographic document processing and further impact a broad spectrum of graphics applications.
Acknowledgements.
The work was partially supported by the following: National Natural Science Foundation of China under No. 92370119, 62436009, 62276258 and 62376113, XJTLU Funding REF-22-01-002.
       (Input)       (Zhang et al., 2022)     (Verhoeven et al., 2023)     (Li et al., 2023a)       (Li et al., 2023b)        (Yu et al., 2024)        (March,25th,2025)        (Our)
Figure 8. More Qualitative comparisons on the DocUNet benchmark. We highlight some obvious content edges with red dotted lines.
Figure 9. More Qualitative comparisons on the AnyPhotoDoc6300 benchmark. We highlight some obvious content edges with red dotted lines.
Figure 11. More Quantitative dewarping performance comparisons on the AnyPhotoDoc6600 benchmark dataset. Under the fixed ”Single-column Paper” Layout Pattern, we provided 18 dimensions of evaluation for AD, MS-SSIM, and MMCER.
Figure 12. More Quantitative dewarping performance comparisons on the AnyPhotoDoc6600 benchmark dataset. Under the fixed ”Magazine” Layout Pattern, we provided 18 dimensions of evaluation for AD, MS-SSIM, and MMCER.